# Brilla AI: AI Contestant for the National Science and Maths Quiz

### v1

@@ 1 Introduction
Motivated by this need, Boateng et al. proposed a grand challenge in education — NSMQ AI Grand Challenge — which is a robust, real-world challenge in education for such an AI: “Build an AI to compete in Ghana’s National Science and Maths Quiz Ghana (NSMQ) competition and win — performing better than the best contestants in all rounds and stages of the competition[2]. In that work, they detailed the motivation for the challenge along with key technical challenges that have to be addressed for an AI to win the NSMQ. We created an open-source project  [^1] Brilla AI Open-Source Project: https://github.com/nsmq-ai/nsmqai to conquer this grand challenge and built Brilla AI as the first key output. Brilla AI is an AI contestant that we deployed to unofficially compete remotely and live in the Riddles round of the 2023 NSMQ Grand Finale, the first of its kind in the 30-year history of the competition.
@@
The NSMQ is an exciting, annual live science and mathematics quiz competition for senior secondary school students in Ghana in which 3 teams of 2 students compete by answering questions across biology, chemistry, physics, and math in 5 rounds over 5 progressive stages until a winning team is crowned for that year [13]. The competition has been run for 30 years and poses interesting technical challenges across speech-to-text, text-to-speech, question-answering, and human-computer interaction. Given the complexity of the challenge, we decided to start with one round, round 5 — Riddles. This round, the final one, is arguably the most exciting as the winner of the competition is generally determined by the performance in the round. In the Riddles round, students answer riddles across Biology, Chemistry, Physics, and Mathematics. Three (3) or more clues are read to the teams that compete against each other to be first to provide an answer (usually a word or a phrase) by ringing their bell. The clues start vague and get more specific. To make it more exciting and encourage educated risk-taking, answering on the 1st clue fetches 5 points, on the 2nd clue — 4 points, and on the 3rd or any clue thereafter, 3 points. There are 4 riddles for each contest with each riddle focusing on one of the 4 subjects. Speed and accuracy are key to winning the Riddles round. An example riddle with clues and the answer is as follows (see a live example here:  [^2] Video Example of Riddle: https://www.youtube.com/watch?v=kdaxoFjiYJg. Question: (1) I am a property of a periodic propagating disturbance. (2) Therefore, I am a property of a wave. (3) I describe a relationship that can exist between particle displacement and wave propagation direction in a mechanical wave. (4) I am only applicable to waves for which displacement is perpendicular to the direction of wave propagation. (5) I am that property of an electromagnetic wave which is demonstrated using a polaroid film. Who am I? Answer: Polarization.
@@
Here are some of the key technical challenges an AI system will need to address. How can it accurately provide real-time transcripts of Ghanaian-accented riddle questions read in English? Speech-to-text systems tend to be trained using data from the West and generally do not work well for African-accented speech [14]. How can it infer the start and end of each riddle and extract only the clues while discarding all statements (e.g., the quiz mistress pausing the reading of the clues after the ring of a bell to listen to an answer from a school, instructions that are read at the start of the round, etc.)? How does the AI know the optimal time to attempt to provide an answer optimizing for speed and accuracy? Providing an answer too early might result in a wrong answer given that earlier clues tend to be vague. Waiting too long could result in an answer being provided by one of the contesting teams. How can the AI say its answer with a Ghanaian accent? Similar to STT systems, several text-to-speech tend to be trained using data from the West and generally do not speak with African accents. How do we ensure all these work seamlessly in real-time without any noticeable latency?
@@ 3 Brilla AI System
Brilla AI is currently available as a web app (built with Streamlit) that livestreams the Riddles round of the contest, and runs 4 machine-learning (ML) systems in real-time via FastAPI together running as an AI server on Google Colab: (1) speech-to-text (using Whisper [16]) which transcribes Ghanaian-accented speech of scientific riddles (2) question extraction (using BERT [5]) which extracts relevant portions of the riddles (clues) by inferring the start and end of each riddle and segmenting the clues (3) question answering (using Mistral [8]) which provides an answer to the riddle and (4) text-to-speech (using VITS [10]) which says the answer with a Ghanaian accent (Figure 1). We used Streamlit due to its applicability to ML and Data science use cases, and quick development cycles.
@@
The web app extracts audio for the video and chunks it in 5 seconds using ffmeg to allow for l0w-latency processing, and then sends and receives relevant data from the AI server sequentially. First, the audio chunk is sent to the STT API for processing which extracts the transcript, and passes it to the QE service which detects the riddle start and extracts the clues. The clues along with information about whether it is a new riddle are sent to the web app. It displays the transcript and clues and then sends the clues to the QA API which determines whether to provide an answer (details described later). If an answer is generated, it is returned to the web app which then displays it and sends it to the TTS API. It generates an audio of the answer and sends it to the web app which then plays it. This current sequential operation is a bottleneck within the system which caused some latency issues (discussed later). We are exploring parallelizing operations to remove such bottlenecks from the system.
The web app consists of 2 modes: demo and live  [^3] Brilla AI Demos: https://youtube.com/playlist?list=PLl-Yq3F_E8VJWFLpsnizorhIz8N-OqUqw&si=ddrl7u0DBZcbNt7H. The demo mode showcases the AI functionality outside of the quiz season. Functionality has been added for users to experience the AI using curated text, audio, and video content but also allows them to provide real-time content to test the AI. For example, in the demo mode, the user can provide real-time audio input on the Speech-to-Text page. The web app processes this audio and sends an API call to the Speech-to-Text API server. The server returns the real-time transcript displayed in the web app to the user. Also, users can provide a YouTube link of any NSMQ quiz video to the web app and see the AI process the content in real-time and compete for that video. To demonstrate the AI during active quiz season, the web app utilizes a live mode. In this mode, the user can supply a live stream link from YouTube to the web app and see the AI compete in real-time as well.
@@ 4 Data Collection and Preprocessing
We curated a dataset of NSMQ contests from 2012-2022 containing videos of the contest and corresponding metadata, text form of riddles questions, and open-source science textbooks. First, we created a Google sheet with information about each contest such as the contest date, names of the competing schools, and the marks they acquired at the end of each contest. We additionally provided information for the riddles round such as which school answered each riddle and at which clue, which was used to calculate the points obtained by the best human contestants for each riddle. We then used that information to automatically extract video clips of the contests. Next, we manually generated annotations for the riddle round containing information about the start and end timestamps of each riddle and clue. These were then used to automatically exactly the video and audio clip segments of riddles.
@@
We extracted HTML files from the following open-source Science textbooks on OpenStax — High School Physics, College Physics, College Biology, College Chemistry, and College Algebra — segmenting content into chapters, sections, paragraphs, and passages. We have completed the sheet with information about contests (with some contest information missing for some years as they could not be found online). For information about the performance of contestants in the riddle round, we only have complete information years 2019 and 2020. Though we have complete data for the text version of the riddle questions, we only used riddles for these 2 years for evaluation so we could compare the performance of our system with human performances. Consequently, we used a total of 316 riddles (156 in 2019 and 160 in 2020). Work is ongoing to annotate the timestamps of each of the clues for all riddles and also to completely parse the textbooks. Hence, we did not have a complete set of video and audio clips of riddles for this work.
@@ 5.1 Speech-to-Text
The Speech-To-Text (STT) system  [^4] STT System Demo: https://youtu.be/jn4Kh7fNgGs provides a robust, fast transcription service for Ghanaian-accented English speech containing mathematical and scientific content. We used OpenAI’s Whisper model [16] for speech transcription and word error rate (WER) and latency as our evaluation metrics. We previously evaluated all versions of Whisper with a limited sample of 3 audios (approximately 15 seconds long) from the NSMQ competition which consists of speech with Ghanaian accents, along with their corresponding transcripts [3]. The WERs are quite high, making the case to fine-tune the model with Ghanaian accented speech in the future. None of the models attained the best scores for both metrics, which warranted the need for a trade-off between the model metrics. An ideal model would have the lowest WER and latency values. We selected the Medium (English) model to deploy for use as it had a good trade-off between WER and latency.
@@ 5.3 Question Answering
The Question Answering (QA) system  [^5] QA System Demo: https://youtu.be/VfkxZAdZ2PA takes as input a riddle that consists of a clue or set of clues, and then attempts to provide answers swiftly and accurately ahead of human contestants. We implemented and evaluated two approaches: Extractive QA and Generative QA. The Extractive QA approach involves retrieving contexts relevant to the current set of clues from a semantic search engine such as our custom-built semantic search engine consisting of a custom vector database computed with Sentence-BERT (SBERT) [17] over passages from the Simple English Wikipedia dataset or Kwame for Science which using SBERT and a science dataset + Simple English Wikipedia [1]. We passed the concatenated clues into the semantic search engine and retrieved the top three passages with the highest similarity scores. We then passed the concatenated clues into an extractive QA model (DistilBERT) as “questions”, while the retrieved passages are provided as the “context” to return a span of text as an answer. The Generative QA approach uses generative models (Falcon-7b-Instruct and Mistral-7B-Instruct-v0.1) along with a well-developed prompt to generate an answer given a set of riddle clues. We made use of prompt engineering to develop a high-quality prompt that takes the clues as part of the prompt. Furthermore, we implemented a confidence modeling pipeline that produces an estimate of the confidence of the QA system to decide whether or not to attempt to answer the riddle after a clue has been received or wait for more clues. For a riddle, we pass the first clue as an input to the model, generate three answers to the riddle, and keep a running count of the number of times each answer is generated for subsequent input clues and compare against a threshold value which we determine empirically after evaluating different threshold values. If any answer’s count is equal to or more than the threshold value, we then return that answer as our answering attempt.
@@
We perform two sets of evaluations. The first evaluation uses all clues concatenated as input to our models and assesses performance (Table 1). This case assumes that the QA system has all the clues and enables us to compare different models. The second evaluation uses the best approach and model from the first evaluation and integrates the confidence modeling approach to simulate the live deployment offline (Table 2). Given that the live mode uses an automatic transcript of 5 seconds of audio, we estimated how many words are in such a chunk (about 7 words). We break our text of the riddles into those chunks and pass them one after another while appending subsequent chunks as input to the model as it attempts to answer at some point. We compared the performance of our 2 approaches — Extractive QA (using SBERT + TQA or Simple Wiki for retrieval and DistilBERT for extraction) and Generative QA (using Mistral-7b-Instruct-v0.1 and Falcon-7b-Instruct) with GPT 3.5 and a human benchmark which we computed using our annotations about the number of the clue on which the contestanting students answered the riddles. For all experiments and evaluations, we run on either a Tesla T4 GPU with 16GB VRAM, or a single A100 GPU with 40GB VRAM.
Our results are shown in Tables 1 and 2). From Table 1, our generative approach using Mistral performed better than the extractive approach but worse than GPT 3.5 and students. Hence, then used it for our second evaluation with confidence modeling and deployed it. GPT 3.5 performed better than our approach but worse than the best human contestants for both evaluations 1 and 2 (Table 2). Overall, these show advancements in generative models for solving science questions. However, the proprietary and closed nature of GPT 3.5, underscores the need to create more accurate open alternatives that will be accessible to students and trainers in Africa and other low-resourced environments. Fine-tuning our open-source models in the future could potentially accomplish that.
@@ 5.4 Text-to-Speech
We performed two evaluation experiments using the ONNX versions of the models. Evaluation 1 involved synthesizing 30 samples of scientific and mathematical speech from past NSMQ questions and Evaluation 2 involved synthesizing 30 samples of conversational speech. We compared the 3 speaker-specific models. We used Mean Opinion Score (MOS), Word Error Rate (WER), and latency as evaluation metrics. The automatic MOS is an objective evaluation of how ‘natural’ the synthesized speech sounds with a range from 1 (bad) to 5 (excellent) [12]. The WER of the synthesized speech measures the intelligibility of the synthesized speech, whereas the latency measures the inference speed of the model. The results (Table 3) show that the Speaker 1 model achieves the best WER. The models generally exhibit high MOS scores, indicating a human-like sound. However, they struggle with scientific and mathematical text synthesis as shown by the poor WER. Qualitatively, audios synthesized with these models sound Ghanaian and similar to the original speakers, with a slight robotic undertone and suboptimal intelligibility for single words. We did not use Speaker 1 to avoid confusion in the live deployment as since it is the voice of the quiz mistress. Consequently, we deployed Speaker 3 due to its lower latency and WER compared to Speaker 2.
@@ 6 Real-World Deployment and Evaluation of Brilla AI
In its debut in the NSMQ Grand Finale in October 2023, Brilla AI answered one of the 4 riddles ahead of the 3 human contesting teams, unofficially placing second (tied) and achieving a 25% EM accuracy  [^8] Brilla AI Debut: https://youtu.be/2AUpiVB6zA4 [9]. For the first riddle, the QE component could not detect the start of the riddle as STT wrongly transcribed “first riddle” as “test riddle”. Consequently, no clues were extracted, QA received no data, and thus could not answer the riddle (the answer shown was from a previous test attempt). For the second and third riddles, the start and ends were detected and the clues were extracted but QA attempted too early (perhaps due to the confidence threshold being quite low) and got the answer wrong. For the fourth and final riddle, QE detected the start of the riddle, could not extract clue one, but extracted the second clue and third clue, and then QA generated an answer, resulting in Brilla AI answering correctly before one of the students answered correctly also (which he did after all the clues). Overall, we recorded a success rate of 25% unofficially placing second (tied) as the best-performing team answered 2 riddles correctly (excluding the one the AI answered), one team answered only one and the third answered none. This result is an important milestone, especially given it was our first live deployment and there were a lot of real-world challenges like noise, inaccurate transcripts, undetected clues, and irrelevant data sent along to QA. Our automatic transcripts were not always accurate and some clues were not detected which all compounded and posed challenges for QA. Our real-time transcription began to lag the live stream of the contest because of the sequential processing of our pipeline, which resulted in our AI appearing to have provided an answer after all clues, even though it did so only after receiving the 3rd clue. This issue could have resulted in a late answer.

### v2

@@ 1 Introduction
Motivated by this need, Boateng et al. proposed a grand challenge in education — NSMQ AI Grand Challenge — which is a robust, real-world challenge in education for such an AI: “Build an AI to compete in Ghana’s National Science and Maths Quiz Ghana (NSMQ) competition and win — performing better than the best contestants in all rounds and stages of the competition[2]. In that work, they detailed the motivation for the challenge and key technical challenges that must be addressed for an AI to win the NSMQ. We created an open-source project  [^1] Brilla AI Open-Source Project: https://github.com/nsmq-ai/nsmqai to conquer this grand challenge and built Brilla AI as the first key output which extends our prior work [3]. Brilla AI is an AI contestant that we deployed to unofficially compete remotely and live in the Riddles round of the 2023 NSMQ Grand Finale, the first of its kind in the 30-year history of the competition.
@@
The NSMQ is an exciting, annual live science and mathematics quiz competition for senior secondary school students in Ghana in which 3 teams of 2 students compete by answering questions across biology, chemistry, physics, and math in 5 rounds over 5 progressive stages until a winning team is crowned for that year [13]. The competition has been run for 30 years and poses interesting technical challenges across speech-to-text, text-to-speech, question-answering, and human-computer interaction. Given the complexity of the challenge, we decided to start with one round, round 5 — Riddles. This round, the final one, is arguably the most exciting as the competition’s winner is generally determined by the performance in the round. In the Riddles round, students answer riddles across Biology, Chemistry, Physics, and Mathematics. Three (3) or more clues are read to the teams that compete against each other to be first to provide an answer (usually a word or a phrase) by ringing their bell. The clues start vague and get more specific. To make it more exciting and encourage educated risk-taking, answering on the 1st clue fetches 5 points, on the 2nd clue — 4 points, and on the 3rd or any clue thereafter, 3 points. There are 4 riddles for each contest with each riddle focusing on one of the 4 subjects. Speed and accuracy are key to winning the Riddles round. An example riddle with clues and the answer is as follows (see a live example here:  [^2] Video Example of Riddle: https://www.youtube.com/watch?v=kdaxoFjiYJg. Question: (1) I am a property of a periodic propagating disturbance. (2) Therefore, I am a property of a wave. (3) I describe a relationship that can exist between particle displacement and wave propagation direction in a mechanical wave. (4) I am only applicable to waves for which displacement is perpendicular to the direction of wave propagation. (5) I am that property of an electromagnetic wave which is demonstrated using a polaroid film. Who am I? Answer: Polarization.
@@
Here are some of the key technical challenges an AI system will need to address. How can it accurately provide real-time transcripts of Ghanaian-accented riddle questions read in English? Speech-to-text systems tend to be trained using data from the West and generally do not work well for African-accented speech [14]. How can it infer the start and end of each riddle and extract only the clues while discarding all statements (e.g., the quiz mistress pausing the reading of the clues after the ring of a bell to listen to an answer from a school, instructions that are read at the start of the round, etc.)? How does the AI know the optimal time to attempt to provide an answer optimizing for speed and accuracy? Providing an answer too early might result in a wrong answer given that earlier clues tend to be vague. Waiting too long could result in an answer being provided by one of the contesting teams. How can the AI say its answer with a Ghanaian accent? Similar to STT systems, several text-to-speech systems tend to be trained using data from the West and generally do not speak with African accents. How do we ensure all these work seamlessly in real time without any noticeable latency?
@@ 3 Brilla AI System
Brilla AI is currently available as a web app (built with Streamlit) (Fig 1) that livestreams the Riddles round of the contest, and runs 4 machine-learning (ML) systems in real-time via FastAPI together running as an AI server on Google Colab: (1) speech-to-text (using Whisper [16]) which transcribes Ghanaian-accented speech of scientific riddles (2) question extraction (using BERT [5]) which extracts relevant portions of the riddles (clues) by inferring the start and end of each riddle and segmenting the clues (3) question answering (using Mistral [8]) which provides an answer to the riddle and (4) text-to-speech (using VITS [10]) which says the answer with a Ghanaian accent (Figure 2). We used Streamlit due to its applicability to ML and Data science use cases, and quick development cycles.
Figure 1: Screenshot of Brilla AI web app
@@
Here is an overview of the sequential processing and communication between components in the Brilla AI system. The web app extracts audio from the video and chunks it in 5 seconds using FFmpeg to allow for low-latency processing, and then sends and receives relevant data from the AI server sequentially. First, the audio chunk is sent to the STT API for processing which extracts the transcript, and passes it to the QE service which detects the riddle start and extracts the clues. The clues along with information about whether it is a new riddle are sent to the web app. It displays the transcript and clues and then sends the clues to the QA API which determines whether to provide an answer (details described later). If an answer is generated, it is returned to the web app which then displays it and sends it to the TTS API. It generates an audio of the answer and sends it to the web app which then plays it. This current sequential operation is a bottleneck within the system which caused some latency issues (discussed later). We are exploring parallelizing operations to remove such bottlenecks from the system.
Figure 3: Sequential Processing in Brilla AI System
The web app consists of 2 modes: demo and live  [^3] Brilla AI Demos: https://youtube.com/playlist?list=PLl-Yq3F_E8VJWFLpsnizorhIz8N-OqUqw&si=ddrl7u0DBZcbNt7H. The demo mode showcases the AI functionality outside of the quiz season. Functionality has been added for users to experience the AI using curated text, audio, and video content but also allows them to provide real-time content to test the AI. For example, in the demo mode, the user can provide real-time audio input on the Speech-to-Text page. The web app processes this audio and sends an API call to the Speech-to-Text API server. The server returns the real-time transcript displayed in the web app to the user. Also, users can provide a YouTube link of any NSMQ quiz video to the web app and see the AI process the content in real time and compete for that video. To demonstrate the AI during active quiz season, the web app utilizes a live mode. In this mode, the user can supply a live stream link from YouTube to the web app and see the AI compete in real-time as well.
@@ 4 Data Collection and Preprocessing
We curated a dataset of NSMQ contests from 2012-2022 containing videos of the contest and corresponding metadata, text form of riddles questions, and open-source science textbooks. First, we created a Google sheet with information about each contest such as the contest date, names of the competing schools, and the marks they acquired at the end of each contest. We additionally provided information for the riddles round such as which school answered each riddle and at which clue, which was used to calculate the points obtained by the best human contestants for each riddle. We then used that information to automatically extract video clips of the contests. Next, we manually generated annotations for the riddle round containing information about the start and end timestamps of each riddle and clue. These were then used to automatically extract the video and audio clip segments of riddles.
@@
We extracted HTML files from the following open-source Science textbooks on OpenStax — High School Physics, College Physics, College Biology, College Chemistry, and College Algebra — segmenting content into chapters, sections, paragraphs, and passages. We have completed the sheet with information about contests (with some contest information missing for some years as they could not be found online). For information about the performance of contestants in the riddle round, we only have complete information for the years 2019 and 2020. Though we have complete data for the text version of the riddle questions, we only used riddles for these 2 years for evaluation so we could compare the performance of our system with human performances. Consequently, we used a total of 316 riddles (156 in 2019 and 160 in 2020). Work is ongoing to annotate the timestamps of each of the clues for all riddles and also to completely parse the textbooks. Hence, we did not have a complete set of video and audio clips of riddles for this work.
@@ 5.1 Speech-to-Text
The Speech-To-Text (STT) system  [^4] STT System Demo: https://youtu.be/jn4Kh7fNgGs provides a robust, fast transcription service for Ghanaian-accented English speech containing mathematical and scientific content. We used OpenAI’s Whisper model [16] for speech transcription and word error rate (WER) and latency as our evaluation metrics. We previously evaluated all versions of Whisper with a limited sample of 3 audios (approximately 15 seconds long) from the NSMQ competition which consists of speech with Ghanaian accents, along with their corresponding transcripts [3]. The WERs are quite high, making the case to fine-tune the model with Ghanaian accented speech in the future. None of the models attained the best scores for both metrics, which warranted the need for a trade-off between the model metrics. An ideal model would have the lowest WER and latency values. We selected the Medium (English) model to deploy for use as it had a good trade-off between WER and latency (31.61% and 0.94 seconds respectively).
@@ 5.3 Question Answering
The Question Answering (QA) system  [^5] QA System Demo: https://youtu.be/VfkxZAdZ2PA takes as input a riddle that consists of a clue or set of clues, and then attempts to provide answers swiftly and accurately ahead of human contestants. We implemented and evaluated two approaches: Extractive QA and Generative QA.
The Extractive QA approach involves retrieving contexts relevant to the current set of clues from a semantic search engine such as our custom-built semantic search engine consisting of a custom vector database computed with Sentence-BERT (SBERT) [17] over passages from the Simple English Wikipedia dataset or Kwame for Science which used SBERT and a science dataset + Simple English Wikipedia [1]. We passed the concatenated clues into the semantic search engine and retrieved the top three passages with the highest similarity scores. We then passed the concatenated clues into an extractive QA model (DistilBERT) as “questions”, while the retrieved passages are provided as the “context” to return a span of text as an answer.
The Generative QA approach uses generative models (Falcon-7b-Instruct and Mistral-7B-Instruct-v0.1) along with a well-developed prompt to generate an answer given a set of riddle clues. We made use of prompt engineering to develop a high-quality prompt that takes the clues as part of the prompt. Furthermore, we implemented a confidence modeling pipeline that produces an estimate of the confidence of the QA system to decide whether or not to attempt to answer the riddle after a clue has been received or wait for more clues. For a riddle, we pass the first clue as an input to the model, generate three answers to the riddle, and keep a running count of the number of times each answer is generated for subsequent input clues and compare against a threshold value which we determine empirically after evaluating different threshold values. If any answer’s count is equal to or more than the threshold value, we then return that answer as our answering attempt.
@@
We perform two sets of evaluations. The first evaluation uses all clues concatenated as input to our models and assesses performance (Table 1). This case assumes that the QA system has all the clues and enables us to compare different models. The second evaluation uses the best approach and model from the first evaluation and integrates the confidence modeling approach to simulate the live deployment offline (Table 2). Given that the live mode uses an automatic transcript of 5 seconds of audio, we estimated how many words are in such a chunk (about 7 words). We break our text of the riddles into those chunks and pass them one after another while appending subsequent chunks as input to the model as it attempts to answer at some point. We compared the performance of our 2 approaches — Extractive QA (using SBERT + Kwame for Science dataset or Simple Wiki for retrieval and DistilBERT for extraction) and Generative QA (using Mistral-7b-Instruct-v0.1 and Falcon-7b-Instruct) with GPT 3.5 and a human benchmark which we computed using our annotations about the clue number on which the contestanting students answered the riddles. For all experiments and evaluations, we run on either a Tesla T4 GPU with 16GB VRAM, or a single A100 GPU with 40GB VRAM on Google Colab.
Our results are shown in Tables 1 and 2. From Table 1, our generative approach using Mistral performed better than the extractive approach but worse than GPT 3.5 and students. Hence, we used Mistral for our second evaluation with confidence modeling and deployed it. GPT 3.5 performed better than our approach but worse than the best human contestants for both evaluations 1 and 2 (Table 2). Overall, these show advancements in generative models for solving science questions. However, the proprietary and closed nature of GPT 3.5, underscores the need to create more accurate open alternatives that will be accessible to students and trainers in Africa and other low-resourced environments. Fine-tuning our open-source models in the future could potentially accomplish that.
@@ 5.4 Text-to-Speech
We performed two evaluations using the ONNX versions of the models. Evaluation 1 involved synthesizing 30 samples of scientific and mathematical speech from past NSMQ questions and Evaluation 2 involved synthesizing 30 samples of conversational speech. We compared the 3 speaker-specific models. We used Mean Opinion Score (MOS), Word Error Rate (WER), and latency as evaluation metrics. The automatic MOS is an objective evaluation of how ‘natural’ the synthesized speech sounds with a range from 1 (bad) to 5 (excellent) [12]. The WER of the synthesized speech measures the intelligibility of the synthesized speech, whereas the latency measures the inference speed of the model. The results (Table 3) show that the Speaker 1 model achieves the best WER. The models generally exhibit high MOS scores, indicating a human-like sound. However, they struggle with scientific and mathematical text synthesis as shown by the poor WER. Qualitatively, audios synthesized with these models sound Ghanaian and similar to the original speakers, with a slight robotic undertone and suboptimal intelligibility for single words. We did not use Speaker 1 to avoid confusion in the live deployment since it is the voice of the quiz mistress. Consequently, we deployed Speaker 3 due to its lower latency and WER compared to Speaker 2.
@@ 6 Real-World Deployment and Evaluation of Brilla AI
In its debut in the NSMQ Grand Finale in October 2023, Brilla AI answered one of the 4 riddles ahead of the 3 human contesting teams, unofficially placing second (tied) and achieved a 25% EM accuracy  [^8] Brilla AI Debut: https://youtu.be/2AUpiVB6zA4 [9]. For the first riddle, the QE component could not detect the start of the riddle as STT wrongly transcribed “first riddle” as “test riddle”. Consequently, no clues were extracted, QA received no data, and thus could not answer the riddle (the answer shown was from a previous test attempt). For the second and third riddles, the start and ends were detected and the clues were extracted but QA attempted too early (perhaps due to the confidence threshold being quite low) and got the answer wrong. For the fourth and final riddle, QE detected the start of the riddle, could not extract clue one, but extracted the second clue and third clue, and then QA generated an answer, resulting in Brilla AI answering correctly before one of the students answered correctly also (which he did after all the clues). Overall, we recorded a success rate of 25% unofficially placing second (tied) as the best-performing team answered 2 riddles correctly (excluding the one the AI answered), one team answered only one and the third answered none. This result is an important milestone, especially given it was our first live deployment and there were a lot of real-world challenges like noise, inaccurate transcripts, undetected clues, and irrelevant data sent along to QA. Our automatic transcripts were not always accurate and some clues were not detected which all compounded and posed challenges for QA. Our real-time transcription began to lag the live stream of the contest because of the sequential processing of our pipeline, which resulted in our AI appearing to have provided an answer after all clues, even though it did so only after receiving the 3rd clue. This issue could have resulted in a late answer.